# Joint identification of spatially variable genes via a network-assisted Bayesian regularization approach

### v1

@@ 3.1 Basic Simulations
For evaluating identification performance, we adopt Recall = $\frac{\text{TP+FN}}{\operatorname{TP}}$ , Precision =$\frac{\operatorname{TP+FP}}{\operatorname{TP}}$, and F1 score = $\frac{2\cdot\text{Precision}\cdot\operatorname{Recall}}{\text{Precision+Recall}},$ with TP, FP, and FN being the numbers of true positives, false positives, and false negatives, respectively. Under each scenario, we simulate 50 replicates. Summary results under the scenarios with a low dropout rate are given in Table 1. The rest of the results with a high dropout rate are provided in Supplementary Table S1. It is observed that the proposed approach achieves superior accuracy in identifying SV genes with higher F1 scores across all scenarios. Overall, CTSV-g achieves the second-best identification performance since it also accommodates dropouts and over-dispersion through the adoption of ZINB distribution. Improved performance over CTSV-g supports the validity of incorporating network information and accommodating cellular diversity. Furthermore, the majority of the alternatives exhibit good performance with the simple linear patterns while the proposed approach performs stably across different spatial expression patterns. Moreover, similar to those in some published studies, SPARK-X exhibits many false negatives, probably due to its covariance test framework and neglection of cellular composition confounding, and SPARK achieves excellent FP control across all scenarios, mainly attributable to the robust Cauchy combination rule. CTSV identifies the fewest global SV genes across all scenarios due to its concern with cell-type-specific SV genes and inability in accurately identifying genes with global spatial expression variations. Noticeably, under the scenarios with a scale-free network where the signals of the SV genes are larger, all approaches exhibit improved performance with higher F1 score values. However, the superiority of the proposed approach is again evidently observed. With a higher dropout rate, which is more common with practical ST data, all approaches have decayed performance, especially for SpatialDE which is not well-suited for sparse expression distribution. However, the superiority of the proposed approach becomes more prominent.
@@ 3.2 Model Misspecification
We further evaluate performance of the proposed approach under scenarios where the data generation model is misspecified. In particular, we consider two types of model misspecification: (M1) the ZINB model considered in Yu and Luo (2022) where the logarithmic mean value is set as a mix of cell-type-specific expression levels; and (M2) the Poisson model considered in Sun et al. (2020) with the parameters inferred from a real mouse olfactory bulb study. More detailed settings are provided in Supplementary Section S3.4, and summary results are reported in Figure S6.
It is observed that the proposed approach can still maintain its superiority or at least perform comparably to the method that matches the data generation model. Specifically, under model (M1), which favors the CTSV method, the proposed approach achieves more accurate identification with a median F1 equal to 1, while CTSV exhibits the second-best performance. Moreover, even under model (M2) without cellular composition confounding, the proposed approach can still correctly identify SV genes with a median F1 score larger than 0.92. Here, spatialDE and SPARK also behave well since the expression pattern is inferred based on the analysis results of the two approaches.
@@ 3.3 Examination on the Noise of Network and Cellular Composition
We continue to examine scenarios with noisy network and cellular composition information. In particular, the noisy network is likely to involve edges that connect both SV and non-SV genes. Two specific generation models are considered, including the ZINB model considered in Section 3.1 with a star network, a linear spatial pattern, and two degrees of sparsity, and the misspecified model (M1) considered in Section 3.2, where in each informative sub-network, about 20% of the genes connected to the TF genes are uninformative. The summary results are reported in Figure S7. It can be seen that the proposed approach can still achieve an effective balance between the prior network information and observation likelihood, leading to certain robustness.
Moreover, to explore the impact of noisy cellular composition on identification performance, we consider the scale-free network with the linear pattern and high sparsity (dropout rate = 0.5), where the cell-type deconvolution estimates $\hat{\boldsymbol{w}}_{i}$ are sampled from $\operatorname{Dirc}(c\cdot\boldsymbol{w}_{i})$ with $\boldsymbol{w}_{i}$ being the true value and concentration parameter $c=100,80,$ and 50. The lower value of $c$ indicates less accuracy of the cellular composition estimates. The summary results are provided in Figure S8. As expected, the noisy cellular compositions tend to result in decreased identification performance with increased variance. However, even with $c=50$, the identification performance is still satisfactory, and the superiority over the alternatives maintains.
@@ 4 Data Analysis
We apply the proposed approach to the triple-negative breast cancer (TNBC) dataset from a cohort study of TNBC tumors (Bassiouni et al., 2023). A total of 28 tissue sections representing 14 primary TNBC tumors are subject to spatial transcriptomics using the 10x Genomics Visium platform. The raw expression data and the tissue information are available at the Gene Expression Omnibus (GEO) under record GSE210616. Here, we focus on the first tissue section from patient ID 1, which contains 1,109 spots and 36,601 genes. Following the published studies (Charitakis et al., 2023; Yan and Luo, 2024), to improve efficiency and stability, we first remove the genes expressed in fewer than 5 spots and the spots with fewer than 500 expressed genes, resulting in a total of 1,108 spots and 17,781 genes, and then select the top 5,000 highly variable genes for downstream analysis. After preprocessing, the median value of expressed counts across spots is 7,036, and the proportion of zero expressions is about 68.89%.
We consider the protein-protein interaction (PPI) network obtained from the STRING database (Szklarczyk et al., 2023). Integrating the PPI network has been widely adopted in recent breast cancer analysis, which provides a powerful complement for a mechanistic understanding of the genomic alterations related to breast cancer (Kim et al., 2021). In particular, it has been demonstrated in published studies (Pranavathiyani et al., 2019; Chen et al., 2021) that the oncogenes or prognostic genes associated with breast cancer are typically highly interconnected within distinct modules, underscoring the importance of considering the connections within networks in SV gene identification. Matching with this PPI network leads to 1,687 connected genes involved in 164 disconnected sub-networks, while the remaining 3,313 genes are singleton nodes.
@@
Figure 2: (A) Pie charts of the distributions of cell types across spots. (B) Spatial expression patterns of three representative SV genes identified by all of SPARK-X, spatialDE, SPARK, HEARTSVG, and MERINGUE. (C) Spatial expression patterns of three representative SV genes only identified by the proposed approach.
@@
The proposed approach identifies 1,084 SV genes. Among them, 782 are connected in the PPI network, of which the graph representation is shown in Figure 3(A). Analysis is also conducted using the alternatives. The upset plot, which provides the numbers of the SV genes identified by different approaches as well as their overlaps, is shown in Figure 4. Specifically, SPARK-X identifies the largest number of SV genes (2,365), while CTSV detects the smallest (266). A total of 54 overlapping genes are identified by all seven approaches. For SPARK-X, spatialDE, SPARK, HEARTSVG, and MERINGUE, there are more overlapping SV genes (245), mainly due to the potential cell-type diversity confounding that is not well accommodated by these approaches. To provide a more intuitive illustration, among these 245 genes, we take three representative genes MICAL2, ADAM12, and LTBP2 and present their spatial expression patterns in Figure 2(B). All three genes are observed to have significant cell type diversity. Moreover, 47 genes (all connected) are only selected by the proposed approach. Among them, the expression patterns of three representative genes SF3B4, PCGF6, and MAPK15 are presented in Figure 2(C), where clear spatial variability (not driven by cell type diversity) is observed.
Figure 3: Network of 782 connected SV genes identified by the proposed approach. (B) Three representative detected communities in the network.
@@
Figure 4: Upset plot of the numbers of SV genes identified by different approaches and their overlaps.
@@
To provide additional comparisons for the SV genes identified by different approaches, we obtain the lists of breast cancer driver genes from the Candidate Cancer Gene Database (CCGD), IntOGen, and DriverDB databases, which contain 579, 290, and 222 genes, respectively, and examine whether these genes are identified as SV. The inclusion rates per thousand are (20.30, 11.70, 23.06) for the proposed approach, (13.26, 7.50, 14.42) for spatialDE, (14.58, 7.61, 13.95) for SPARK, (19.45, 5.50, 14.80) for SPARK-X, (0.00, 0.00, 7.52) for CTSV, (12.37, 9.00, 13.50) for MERINGUE, and (12.44, 8.61, 6.70) for HEARTSVG, respectively. This may support the higher power of the proposed approach to some extent.
@@
We then conduct community detection with the 782 connected SV genes and show three representative communities in Figure 3(B), where the overlapping genes identified by all seven approaches (Overlap), genes only identified by the proposed approach (proposed Only), and the remaining genes (Others) are represented by different colors. It is observed that most overlapping genes are hubs with larger values of degrees. Some of the proposed Only genes are connected to the hub genes, for example, SF3B4 and PINX1 in community B, which may exhibit weak spatial patterns but are involved in the same functional organization or regulation mechanism with the hub ones. Some other genes seem to serve as bridge linkage genes with high betweenness centrality. Examples include SGMS2 in community C, which expedites the formulation of the complete gene regulatory networks.
@@
Then, we conduct Gene Ontology (GO) enrichment analysis. Specifically, for the three communities in Figure 3(B), we list the corresponding top five significant GO terms in Table 2, which have important biological implications for breast cancer. For example, in community A, wound healing has been shown to promote the growth of breast cancer, while integrin-mediated signaling has been confirmed to play a significant role in TNBC bone metastasis. The genes involved in community B are mainly enriched in RNA splicing, and the relationships between RNA splicing and a variety of cancers have long been recognized and widely utilized in cancer diagnosis and therapy. For community C, the sphingolipid metabolism has been shown to be essential for breast cancer progression, while ceramide metabolism balance has been confirmed as a critical step of breast cancer development and has been long adopted as a targeted pathway to induce apoptosis in breast cancer cell lines. References that support the above results are presented in Supplementary Section S4.
@@
We further take a closer look at the proposed Only genes. Specifically, we conduct GO enrichment analysis for each community again, but with the proposed Only genes eliminated. The comparison results are shown in Figure 5(A), where the top ten significant GO terms with the proposed Only genes included are considered. It can be seen that when the proposed Only genes are included, some more significant GO terms are found. Moreover, the incorporation of the proposed Only genes contributes to some newly detected GO terms (shown in Figure 5(B)), such as skeletal muscle contraction, which has been confirmed to be critical in tumor progression and therapy response (Crist et al., 2022). These evidences suggest the importance of the proposed Only genes in the PPI network and their biological implications.
@@
GO:0042060 2.20×$\times 10^{-9}$ wound healing
GO:0031589 2.20×$\times 10^{-9}$ cell-substrate adhesion
GO:0007229 3.04×$\times 10^{-8}$ integrin-mediated signaling pathway
GO:0007160 1.07×$\times 10^{-6}$ cell-matrix adhesion
GO:1902905 1.07×$\times 10^{-6}$ positive regulation of supramolecular fiber organization
Community B
GO:0000375 1.19×$\times 10^{-37}$ RNA splicing, via transesterification reactions
GO:0008380 4.22×$\times 10^{-36}$ RNA splicing
GO:0000377 4.22×$\times 10^{-36}$
@@
GO:0000398 4.22×$\times 10^{-36}$ mRNA splicing, via spliceosome
GO:0000387 8.16×$\times 10^{-15}$ spliceosomal snRNP assembly
@@
GO:0030148 1.18×$\times 10^{-14}$ sphingolipid biosynthetic process
GO:0046467 7.13×$\times 10^{-14}$ membrane lipid biosynthetic process
GO:0006665 9.42×$\times 10^{-14}$ sphingolipid metabolic process
GO:0006643 5.62×$\times 10^{-13}$ membrane lipid metabolic process
GO:0006672 9.39×$\times 10^{-13}$ ceramide metabolic process
Figure 5: Lollipop plots of GO enrichment analysis for the three communities. Count represents the number of the input genes that are annotated to the specific GO term, and GeneRatio denotes the proportion of the Count divided by the total number of input genes. (A) Comparison results of the analysis without and with the proposed Only genes. (B) Newly detected significant GO terms with the proposed Only genes included.

### v2

@@ 3.1 Basic Simulations
For evaluating identification performance, we adopt Recall = $\frac{\operatorname{TP}}{\operatorname{TP+FN}}$ , Precision $\frac{\operatorname{TP}}{\operatorname{TP+FP}}$, and F1 score = $\frac{2\cdot\text{Precision}\cdot\operatorname{Recall}}{\text{Precision+Recall}},$ with TP, FP, and FN being the numbers of true positives, false positives, and false negatives, respectively. Under each simulation scenario, we simulate 50 replicates. Summary results under the scenarios with a low dropout rate are given in Table 1. The rest of the results with a high dropout rate are provided in Supplementary Table S2 (Wu et al., 2025). It is observed that the proposed approach achieves superior accuracy in identifying SV genes with higher F1 scores across all scenarios. Overall, CTSV-g achieves the second-best identification performance since it also accommodates dropouts and over-dispersion through the adoption of ZINB distribution. Improved performance over CTSV-g supports the validity of incorporating network information and accommodating cellular diversity. Furthermore, the majority of the alternatives exhibit good performance with the simple linear patterns while the proposed approach performs stably across different spatial expression patterns. Moreover, similar to those in some published studies, SPARK-X exhibits many false positives, probably due to its covariance test framework and neglection of cellular composition confounding, and SPARK achieves excellent FP control across all scenarios, mainly attributable to the robust Cauchy combination rule. HRG also behaves poorly in FP control, which may be attributed to the potential error introduced by low-dimensional projection in the step of graph construction. The other two graph-based approaches, MERINGUE and nnSVG, are more conservative and tend to select fewer SV genes. As a result, they are poor at finding the true signals but show high precision values. CTSV identifies the fewest global SV genes across all scenarios due to its concern with cell-type-specific SV genes and inability in accurately identifying genes with global spatial expression variations. Noticeably, under the scenarios with a scale-free network where the signals of the SV genes are larger, all approaches exhibit improved performance with higher F1 score values. However, the superiority of the proposed approach is again evidently observed. With a higher dropout rate, which is more common with practical ST data, all approaches have decayed performance, especially for SpatialDE which is not well-suited for sparse expression distribution. However, the superiority of the proposed approach becomes more prominent.
3.2 False Positive Rate Assessment Using SV-free Simulated Datasets
In addition to the basic simulations, we conduct another two SV-free simulation scenarios to examine false positive (FP) control performance. Specifically, we consider two different sparsity settings with dropout rates of 0.1 and 0.5. In each setting, we simulate $p=5,000$ non-SV genes while keeping the other parameters the same as those in the above basic simulations.
In the SV-free simulations, for the proposed approach, we once again consider two networks, Star and Scale-free, to examine the FP control performance under different assigned network dependency structures. Comparison boxplots of the False Positive Rate (FPR), calculated as $\frac{\operatorname{FP}}{\operatorname{TN}+\operatorname{FP}}$, based on 50 replicates, are shown in Supplementary Figure S3 (Wu et al., 2025). We omit the results of SINFONIA since it requires a predefined number of target SV genes for identification. We observe similar results to those in the basic simulation. When the dropout rate is lower, SPARK-X, HRG, and HEARTSVG perform poorly in false positive control, with much higher FPRs. In contrast, the remaining approaches all demonstrate satisfactory performance. When the dropout rate is higher, SPARK-X and HEARTSVG tend to select fewer genes in sparse data, resulting in improved false positive control. Meanwhile, the other approaches, except for HRG, can maintain their ability to control false positives.
3.3 Validation for Zero-Inflated Distribution
To assess the necessity of adopting a zero-inflated distribution, we further compare our ZINB model with an NB model (which lacks zero inflation) under varying levels of data sparsity. Specifically, we focus on scenarios involving the star network and periodic spatial patterns, and we present the comparison results in Supplementary Figure S4 (Wu et al., 2025). It is evident that the NB model consistently exhibits reduced performance in SV gene detection. The Recall and F1 metrics decay more significantly as the dropout rate increases. This suggests that zero-inflation modeling aids in recovering spatial signals that might otherwise be obscured by an excess of zeros, particularly under high-sparsity conditions.
3.4 Validation for the Network Assistance Strategy
@@ 3.5 Model Misspecification
We further evaluate performance of the proposed approach under scenarios where the data generation model is misspecified. In particular, we consider three types of model misspecification: (M1) the ZINB model considered in Yu and Luo (2022) where the logarithmic mean value is set as a mix of cell-type-specific expression levels; (M2) the generalized linear spatial model where the spatial variability is introduced through the Gaussian covariance matrix as adopted in Svensson, Teichmann and Stegle (2018) and Sun, Zhu and Zhou (2020); and (M3) the Poisson generalized model considered in Sun, Zhu and Zhou (2020) with the parameters inferred from a real mouse olfactory bulb study. More detailed settings are provided in Supplementary Section S4 (Wu et al., 2025), and summary results are reported in Figure S8.
It is evident that the proposed approach can either maintain its superiority or perform at least as well as the method that aligns with the data generation model. Specifically, under model (M1), which favors the CTSV method, the proposed approach achieves more accurate identification, with a median F1 score of 1. In contrast, CTSV shows the second-best performance. Under model (M2), which involves covariance-based spatial variability, the proposed approach still attains the best F1 performance among all alternative methods, achieving a better balance between Recall and Precision. This enhanced performance once again validates the network assistance strategy. Furthermore, even under model (M3), where there is no cellular composition confounding, the proposed approach can still accurately identify SV genes, with a median F1 score greater than 0.92. In this scenario, spatialDE and SPARK also perform well, as the expression pattern is inferred based on the analysis results of these two approaches.
@@ 3.6 Examination on the Noise of Network and Cellular Composition
We continue to examine scenarios with noisy network and cellular composition information. In particular, the noisy network is likely to involve edges that connect both SV and non-SV genes. Two specific generation models are considered, including the ZINB model considered in Section 3.1 with a star network, a linear spatial pattern, and two degrees of sparsity, and the misspecified model (M1) considered in Section 3.5, where in each informative sub-network, about 20% of the genes connected to the TF genes are uninformative. The summary results are reported in Supplementary Figure S9 (Wu et al., 2025). It can be seen that the proposed approach can still achieve an effective balance between the prior network information and observation likelihood, leading to certain robustness. Moreover, to explore the impact of noisy cellular composition on identification performance, we consider the scale-free network with the linear pattern and high sparsity (dropout rate = 0.5), where the cell-type deconvolution estimates $\hat{\boldsymbol{w}}_{i}$ are sampled from $\operatorname{Dirc}(c\cdot\boldsymbol{w}_{i})$ with $\boldsymbol{w}_{i}$ being the true value and concentration parameter $c=100,80,$ and 50. The lower value of $c$ indicates less accuracy of the cellular composition estimates. The summary results are provided in Figure S10. As expected, the noisy cellular compositions tend to result in decreased identification performance with increased variance. However, even with $c=50$, the identification performance is still satisfactory, and the superiority over the alternatives maintains.
@@ 4 Data Analysis
[…]
The TNBC dataset is obtained from a cohort study of TNBC tumors (Bassiouni et al., 2023), where a total of 28 tissue sections representing 14 primary TNBC tumors are subject to spatial transcriptomics using the 10x Genomics Visium platform. The raw expression data and the tissue information are available at the Gene Expression Omnibus (GEO) under record GSE210616. Here, we focus on the first tissue section from patient ID 1, which contains 1,109 spots and 36,601 genes. Following the published studies (Charitakis et al., 2023; Yan and Luo, 2024), to improve efficiency and stability, we first remove the genes expressed in fewer than 5 spots and the spots with fewer than 500 expressed genes, resulting in a total of 1,108 spots and 17,781 genes, and then select the top 5,000 highly variable genes for downstream analysis. After preprocessing, the median value of expressed counts across spots is 7,036, and the proportion of zero expressions is about 68.89%.
We consider the protein-protein interaction (PPI) network obtained from the STRING database (Szklarczyk et al., 2023). Integrating the PPI network has been widely adopted in recent breast cancer analysis, which provides a powerful complement for a mechanistic understanding of the genomic alterations related to breast cancer (Kim et al., 2021). In particular, it has been demonstrated in published studies (Pranavathiyani et al., 2019; Chen, Verbeek and Wolstencroft, 2021) that the oncogenes or prognostic genes associated with breast cancer are typically highly interconnected within distinct modules, underscoring the importance of considering the connections within networks in SV gene identification. Matching with this PPI network leads to 1,687 connected genes involved in 164 disconnected sub-networks, while the remaining 3,313 genes are singleton nodes.
@@
The proposed approach identifies 1,084 SV genes. Analysis is also conducted using the alternatives. The upset plot, which provides the numbers of the SV genes identified by different approaches as well as their overlaps, is shown in Figure 3. Specifically, SPARK-X identifies the largest number of SV genes (2,365), while CTSV detects the smallest (266). A total of 50 overlapping genes are identified by all 10 approaches. For SPARK-X, SINFONIA, nnSVG, spatialDE, SPARK, HEARTSVG, MERINGUE, and HRG, there are more overlapping SV genes (204), mainly due to the potential cell-type diversity confounding that is not well accommodated by these approaches. To provide a more intuitive illustration, among these 204 genes, we take three representative genes MICAL2, ADAM12, and LTBP2 and present their spatial expression patterns in Figure 4(B). All three genes are observed to have significant cell type diversity. Moreover, among 1,084 genes identified by the proposed approach, 782 are connected in the PPI network, of which the graph representation is shown in Figure 5(A). 47 genes (all connected) are only selected by the proposed approach. Among them, the expression patterns of three representative genes SF3B4, PCGF6, and MAPK15 are presented in Figure 4(C), where clear spatial variability (not driven by cell type diversity) is observed.
@@
Figure 3: Upset plot of the numbers of SV genes identified by different approaches and their overlaps for Visium TNBC data.
Figure 4: (A) Pie charts of the distributions of cell types across spots for Visium TNBC data. (B) Spatial expression patterns of three representative SV genes identified by all of SPARK-X, SINFONIA, nnSVG, spatialDE, SPARK, HEARTSVG, MERINGUE, and HRG. (C) Spatial expression patterns of three representative SV genes only identified by the proposed approach.
Figure 5: (A) Network of 782 connected SV genes identified by the proposed approach for Visium TNBC data. (B) Network of three representative detected communities.
@@
To provide additional comparisons for the SV genes identified by different approaches, we obtain the lists of breast cancer driver genes from the Candidate Cancer Gene Database (CCGD), IntOGen, and DriverDB databases, which contain 579, 290, and 222 genes, respectively, and examine whether these genes are identified as SV. The inclusion rates per thousand are (20.30, 11.70, 23.06) for the proposed approach, (13.26, 7.50, 14.42) for spatialDE, (14.58, 7.61, 13.95) for SPARK, (19.45, 5.50, 14.80) for SPARK-X, (8.39, 12.59, 15.38) for HRG, (0.00, 0.00, 7.52) for CTSV, (12.37, 9.00, 13.50) for MERINGUE, (13.31, 5.58, 13.31) for SINFONIA, (18.62, 6.21, 16.37) for nnSVG, and (12.44, 8.61, 6.70) for HEARTSVG, respectively. This may support the higher power of the proposed approach to some extent.
@@
We then conduct community detection with the 782 connected SV genes and show three representative communities in Figure 5(B), where the overlapping genes identified by all ten approaches (Overlap), genes only identified by the proposed approach (proposed Only), and the remaining genes (Others) are represented by different colors. It is observed that most overlapping genes are hubs with larger values of degrees. Some of the proposed Only genes are connected to the hub genes, for example, SF3B4 and PINX1 in community A, which may exhibit weak spatial patterns but are involved in the same functional organization or regulation mechanism with the hub ones. Some other genes seem to serve as bridge linkage genes with high betweenness centrality. Examples include SGMS1 in community B and STAM2 in community C, which expedites the formulation of the complete gene regulatory networks.
@@
Then, we conduct Gene Ontology (GO) enrichment analysis. Specifically, for the aforementioned three communities, we list the corresponding top five significant GO terms in Table 2, which have important biological implications for breast cancer. For example, the genes involved in community A are mainly enriched in RNA splicing, and the relationships between RNA splicing and a variety of cancers have long been recognized and widely utilized in cancer diagnosis and therapy (Stanley and Abdel-Wahab, 2022; Yamauchi, Nishimura and Yoshimi, 2022). For community B, sphingolipid metabolism has been shown to be essential for breast cancer progression (Corsetto et al., 2023), while the balance of ceramide metabolism has been confirmed as a critical step in breast cancer development and has been long adopted as a targeted pathway to induce apoptosis in breast cancer cell lines (Vethakanraj et al., 2015). In community C, cell cycle phase transition has been confirmed to play a critical role in breast cancer progression, as supported by previous studies (Kashyap et al., 2021).
@@
We further take a closer look at the proposed Only genes. Specifically, we conduct an GO enrichment analysis for each community again, but with the proposed Only genes eliminated. The comparison results are shown in Figure 6 (Left), where the top ten significant GO terms with the proposed Only genes included are considered. It can be seen that when the proposed Only genes are included, some more significant GO terms are found. Moreover, the incorporation of the proposed Only genes contributes to some newly detected GO terms (shown in Figure 6, Right), such as regulation of mRNA metabolic process, which has been confirmed to be critical in targeted therapy, chemotherapy and immunotherapy in breast cancer (Xu et al., 2024). These evidences suggest the importance of the proposed Only genes in the PPI network and their biological implications.
@@
[…]
@@
[…]
Community B
[…]
@@
[…]
Figure 6: Lollipop plots of GO enrichment analysis for the three communities detected in Visium TNBC data. Count represents the number of the input genes that are annotated to the specific GO term, and GeneRatio denotes the proportion of the Count divided by the total number of input genes. Left: Comparison results of the analysis without and with the inclusion of proposed Only genes. Right: Newly detected significant GO terms with the proposed Only genes included.